OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)? -3/2 -2/3 3/2 2/3

OpenStudy (raden):

use formula : slope (m) = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

so the anawer is Y=-3/2

OpenStudy (raden):

m=-6/4=-3/2 y're right

OpenStudy (anonymous):

thanks for you assistance

OpenStudy (raden):

yw